\begin{thebibliography}{10}

% \bibitem{BL}%(MR1727673)
% Y.~Benyamini and J.~Lindenstrauss,
% \newblock {\em Geometric nonlinear functional analysis. {V}ol. 1}, volume~48 of
%   {\em American Mathematical Society Colloquium Publications},
% \newblock American Mathematical Society, Providence, RI, 2000.

% \bibitem{ANP}
% T.~Austin and A.~Naor and Y.~Peres,
% \newblock The wreath product of $\mathbb{Z}$ with $\mathbb{Z}$ has Hilbert compression exponent $\tfrac{2}{3}$,
% \newblock \emph{Proc. Amer. Math. Soc.} \textbf{137}(1):85--90, 2009. 

\bibitem{Baumslag-wrpasfinpres}
G.~Baumslag,
\newblock Wreath products of finitely presented groups,
\newblock \emph{Math. Z.} \textbf{75}:22--28, 1961.

% \bibitem{BS}%()
% I.~Benjamini and O.~Schramm,
% \newblock Every graph with a positive Cheeger constant contains a tree with a positive Cheeger constant, 
% \newblock \emph{Geom. Func. Anal.}, \textbf{7}(3):403--419, 1997.

\bibitem{BP}%()
M.~Bourdon and H.~Pajot,
\newblock Cohomologie $\ell^p$ et espaces de Besov, 
\newblock \emph{J. reine angew. Math.}, \textbf{558}:85--108, 2003.

\bibitem{BCD}
N.~Brady and M.~Clay and P.~Dani,
\newblock Morse theory and conjugacy classes of finite subgroups,
\newblock \emph{Geom. dedicata}, \textbf{147}(1):1--14, 2010.

\bibitem{CG}%(MR837621)
J.~Cheeger and M.~Gromov,
\newblock {$L\sb 2$}-cohomology and group cohomology,
\newblock \emph{Topology}, \textbf{25}:189--215, 1986.

% \bibitem{CKTD}
% Ultraproducts of measure preserving actions and graph combinatorics Clinton T. Conley, Alexander S. Kechris, and Robin D. Tucker-Drob

% \bibitem{Eck}%(MR1760592)
% B.~Eckmann,
% \newblock Introduction to $\ell_2$ methods in topology: reduced $\ell_2$ cohomology, harmonic chains and $\ell_2$ Betti numbers,
% \newblock (notes prepared by Guido Mislin) \emph{Israel J. Math.} \textbf{117}:183--219, 2000. 

\bibitem{El-qi}
G.~Élek,
\newblock Coarse cohomology and $\ell_p$-cohomology,
\newblock $K$-\emph{Theory}, \textbf{13}:1--22, 1998.

\bibitem{Ers}
A.~Erschler,
\newblock On isoperimetric profiles of finitely generated groups,
\newblock \emph{Geom. Dedic.} \textbf{100}(1):157--171, 2003.

% \bibitem{Gab}%(MR2241189)
% D.~Gaboriau,
% \newblock \emph{Invariants $\ell^2$ de relations d'\'equivalence et de groupes},
% \newblock Publ. Math. Inst. Hautes \'Etudes Sci., \textbf{95}:93--150, 2002. 

\bibitem{moi-poiss}
A.~Gournay,
\newblock Boundary values of random walks and $\ell^p$-cohomology in degree one,
\newblock Ar$\chi$i$\nu$:1303.4091

\bibitem{Gro-polyn}
M.~Gromov,
\newblock Groups of Polynomial growth and Expanding Maps.
\newblock \emph{Publications mathematiques I.H.É.S.}, {\bf 53}:53--73 , 1981.

\bibitem{Gro}
M.~Gromov,
\newblock Asymptotic invariants of groups, \textit{in} \emph{Geometric group theory ({V}ol. 2)},
\newblock London Mathematical Society Lecture Note Series, Vol. {\bf 182}, Cambridge University Press, 1993, viii+295.

% \bibitem{Guich}
% A.~Guichardet,
% \newblock Sur la cohomologie des groupes topologiques II. 
% \newblock \emph{Bull. Sci. Math.}, \textbf{96}:305--332, 1972.

% \bibitem{Hall}
% P.~Hall,
% \newblock Finiteness condition for soluble groups. 
% \newblock \emph{Proc. London Math. Soc. (3)}, \textbf{4}:419--436, 1954.

% \bibitem{Hay}
% B.~Hayes, 
% \newblock \emph{An $\ell^p$-version of von Neumann dimension for Banach space representations of sofic groups}
% \newblock Ar$\chi$i$\nu$:1110.5390v3

\bibitem{HS}
I.~Holopainen and P.~Soardi,
\newblock $p$-harmonic functions on graphs and manifolds.
\newblock \emph{Manuscripta Math.}, \textbf{94}, 95--110, 1997.

\bibitem{HS-polyn}
I.~Holopainen and P.~Soardi,
\newblock A strong Liouville theorem for $p$-harmonic functions on graphs.
\newblock \emph{Ann. Acad. Sci. Fenn. Math.}, \textbf{22}, 205--226, 1997.


\bibitem{Kap}
E.~Kappos,
\newblock $\ell^p$-cohomology for groups of type $\FP_n$.
\newblock ar$\chi$i$\nu$:math/0511002 (v2), 2006. 

\bibitem{Kaimano}
V.~A.~Kaimanovich,
\newblock Poisson boundaries of random walks on discrete solvable groups,
\newblock \emph{Probability measures on groups, X}, (Oberwolfach, 1990):205--238, Plenum, New York, 1991. 

% \bibitem{Kanai} 
% M.~Kanai,
% \newblock Rough isometries and combinatorial approximations of geometries of non-compact Riemannian manifolds, 
% \newblock \emph{J. Math. Soc, Japan} \textbf{37}:391--413, 1985.

% \bibitem{Loh}
% N.~Lohoué, 
% \newblock Remarques sur un théorème de Strichartz. 
% \newblock \emph{C. R. Acad. Sci., Ser.I} \textbf{311}:507--510, 1990.

\bibitem{Luck}
W.~Lück,
\newblock \emph{$L^2$-invariants: theory and applications to geometry and $K$-theory},
\newblock Ergebnisse der Mathematik und ihrer Grenzgebiete. 3. Folge., 44. 
\newblock Springer-Verlag, Berlin, 2002. % xvi+595 pp.

% \bibitem{Luck2}
% W.~Lück,
% \newblock $L^2$-invariants from the algebraic point of view \emph{in Geometric and cohomological methods in group theory}, 63--161,
% \newblock London Math. Soc. Lecture Note Ser., 358. Cambridge Univ. Press, Cambridge, 2009.

\bibitem{MV}%(MR2294249)
F.~Martin and A.~Valette,
\newblock On the first $\LL^p$ cohomology of discrete groups.
\newblock \emph{Groups Geom. Dyn.}, \textbf{1}:81--100, 2007.

% \bibitem{MooV}
% S.~Moon and A.~Valette
% \newblock Non-properness of amenable actions on graphs with infinitely many ends. 
% \newblock \emph{Ischia group theory 2006}:227--233. World Sci. Publ., Hackensack, NJ, 2007. 

% \bibitem{OW}%(MR910005)
% D.~S. Ornstein and B.~Weiss,
% \newblock Entropy and isomorphism theorems for actions of amenable groups,
% \newblock \emph{J. Analyse Math.}, \textbf{48}:1--141, 1987.

% \bibitem{Pan}
% P.~Pansu,
% \newblock Cohomologie $L^p$ en degré $1$ des espaces homogènes.
% \newblock \emph{Potential Anal.}, \textbf{27}(2):151--165, 2007.

% \bibitem{Pan-rs}
% P.~Pansu,
% \newblock Cohomologie $L^p$, espaces homogènes et pincement.
% \newblock \emph{Comment. Math. Helv.}, \textbf{83}(2):327--357, 2008. 

\bibitem{Pan-qi}
P.~Pansu,
\newblock Cohomologie $\ell^p$: invariance sous quasi-isométrie.
\newblock Unpublished, but available on P.~Pansu's webpage \texttt{http://www.math.u-psud.fr/\textasciitilde{}pansu/liste-prepub.html}, 1995 (updated in 2004).

\bibitem{amen}
A.~L.~T.~Paterson,
\newblock \emph{Amenability}, Mathematical Surveys and Monographs, \textbf{29}. 
\newblock American Mathematical Society, Providence, RI, 1988. %xx+452 pp. 

\bibitem{amen2}
J.~P.~Pier,
\newblock \emph{Amenable locally compact groups}, Pure and Applied Mathematics (New York). 
\newblock Wiley-Interscience Publication. John Wiley \& Sons, Inc., New York, 1984. %x+418 pp.

\bibitem{Puls-prem}
M.~J.~Puls,
\newblock Group cohomology and $\LL^p$-cohomology of finitely generated groups,
\newblock \emph{Canad. Math. Bull.}, \textbf{46}(2):268--276, 2003. 

\bibitem{Puls-harm}
M.~J.~Puls,
\newblock The first $\LL^p$-cohomology of some finitely generated groups and $p$-harmonic functions,
\newblock \emph{J. Funct. Anal.}, \textbf{237}(2):391--401, 2006. 

\bibitem{Puls-pharmbnd}
M.~J.~Puls,
\newblock Graphs of bounded degree and the $p$-harmonic boundary,
\newblock \emph{Pac. J. Math.}, \textbf{248}(2):429--452, 2010. 

\bibitem{Rud}%(MR1157815)
W.~Rudin,
\newblock ``Functional analysis''. International Series in Pure and Applied Mathematics. Second edition.
\newblock McGraw-Hill, Inc., New York, 1991.

% \bibitem{Sew}
% B.~Seward,
% \newblock  Burnside's Problem, spanning trees, and tilings. 
% \newblock  ar$\chi$i$\nu$:1104.1231 (v2), 2011. 

% \bibitem{Sha}
% M.~Shapiro,
% \newblock Pascal's triangle in Abelian and Hyperbolic groups.
% \newblock \emph{J. Aurstral. Math. Soc. (Series A)}, \textbf{63}:281--288, 1997. 

% \bibitem{Strich}
% R.~Strichartz,
% \newblock Analysis of the Laplacian on a complete Riemannian manifold. 
% \newblock \emph{J. Funct. Anal.} \textbf{52}:48--79, 1983.

\bibitem{Tes}
R.~Tessera,
\newblock Vanishing of the first reduced cohomology with values in an $\LL^p$-representation.
\newblock \emph{Ann. Inst. Fourier}, \textbf{59}(2):851--876, 2009.

\bibitem{Tes-iso}
R.~Tessera,
\newblock Isoperimetric profile and random walks on locally compact solvable groups,
\newblock to appear in \emph{Rev. Mat. Iberoam.}, see also ar$\chi$i$\nu$:0706.4257. 

% \bibitem{Var}
% N.~Varopoulos,
% \newblock Isoperimetric inequalities and Markov chains.
% \newblock \emph{J. Funct. Anal.} \textbf{63}:215--239, 1985.


% \bibitem{Tes-cf}
% R.~Tessera
% \newblock Asymptotic isoperimetry on groups and uniform embeddings into Banach spaces.
% \newblock , 2008.

% \bibitem{Wall}
% C.~Wall
% \newblock Finiteness condition for CW-complexes. 
% \newblock \emph{Ann. Math.}, \textbf{81}:56--69, 1965.

% \bibitem{War-ade}
% A.~Warshall,
% \newblock The asymptotic density of dead ends in non-amenable groups.
% \newblock ar$\chi$i$\nu$:1006.3948 (v1), 2010.

% \bibitem{War-intro}
% A.~Warshall,
% \newblock A group with deep pockets for all finite generating sets.
% \newblock \emph{Israel J. Math.}, \textbf{185}:317--342, 2011. 

% \bibitem{Woe}
% W.~Woess,
% \newblock \emph{Random Walks on Infinite Graphs and Groups}, Cambridge tracts in mathematics, \textbf{138}. 
% \newblock Cambridge University Press, 2000. %xx+452 pp. 

\bibitem{Yam}
M.~Yamasaki,
\newblock Parabolic and hyperbolic infinite networks.
\newblock \emph{Hiroshima Math. J.}, \textbf{7}:135--146, 1977.

\end{thebibliography}

\documentclass[a4paper,11pt,twoside,leqno]{article}
\usepackage{amsmath,amssymb,amsfonts,amsthm,graphicx,fancyhdr}
\usepackage[latin1]{inputenc}
\usepackage[T1]{fontenc}
\usepackage{rotating}
\usepackage[colorlinks=true]{hyperref}

\newtheorem{teo}{Theorem}[section]
\newtheorem{lem}[teo]{Lemma}

\newtheorem{dfn}[teo]{Definition}
\newtheorem{ques}[teo]{Question}

\newtheorem*{tea}{Theorem}

 \newtheoremstyle{drem}% name
      {3pt}%      Space above
      {3pt}%      Space below
      {\rmfamily}%Body font
      {}%         Indent amount (empty = no indent, \parindent = para indent)
      {\bf}%Thm head font
      {:}%        Punctuation after thm head
      {.5em}%     Space after thm head: " " = normal interword space
      {}%         Thm head spec (can be left empty, meaning `normal')

 \theoremstyle{drem}

\newtheorem{exa}[teo]{Example}

\newcommand{\eg}[0]{\emph{e.g.} }
\newcommand{\ie}[0]{\emph{i.e.} }

\newcommand{\ssl}[1]{\underline{#1}}
\newcommand{\srl}[1]{\overline{#1}}

\newcommand{\pgen}[1]{\langle #1 \rangle}

\newcommand{\imp}[0]{\Rightarrow}

\newcommand{\eps}[0]{\varepsilon}

\newcommand{\kk}[0]{\ensuremath{\mathbb{K}}}
\newcommand{\rr}[0]{\ensuremath{\mathbb{R}}}
\newcommand{\zz}[0]{\ensuremath{\mathbb{Z}}}
\newcommand{\nn}[0]{\ensuremath{\mathbb{N}}}

\newcommand{\dd}[0]{\ensuremath{\mathrm{d}\!}}

\newcommand{\Id}[0]{\mathrm{Id}}

\newcommand{\un}[0]{1 \!\! \mathrm{l}} %\mathrm{I} }

\newcommand{\limm}[1]{\textrm{\raisebox{.5ex}{\mbox{$\underset{#1}{\lim}$}}} \:}

\newcommand{\inff}[1]{\textrm{\raisebox{.5ex}{\mbox{$\underset{#1}{\inf}$}}} \:}

\newcommand{\tosp}[1]{\overset{#1}{\to}}

\newcommand{\cay}[0]{\mathrm{Cay}\!}
\newcommand{\diam}[0]{\mathrm{diam}\!}

\newcommand{\trc}[0]{\mathrm{T}\!\mathrm{C}}

\newcommand{\FP}[0]{\mathrm{F}\!\mathrm{P}\!}

\newcommand{\BD}[0]{\mathsf{B}\!\mathsf{D}}
\newcommand{\D}[0]{\mathsf{D}}
\newcommand{\LL}[0]{\mathsf{L}}

\newcommand{\maza}[0]{\mu}

\newcommand{\Gver}[0]{X}
\newcommand{\Grou}[0]{G}

\newcommand{\tpoiss}[0]{\cite[Theorem 1.1]{moi-poiss}}

\begin{document}
\centerline{\Large Vanishing of $\ell^p$-cohomology and transportation cost\footnote{MSC: Primary 20J06; Secondary: 20E22, 31C05, 43A07, 43A15}}

\vspace*{1cm}

\centerline{\large Antoine Gournay\footnote{Université de Neuchâtel, Rue É.-Argand 11, 2000 Neuchâtel, Suisse.}} 
% \textsf{antoine.gournay@unine.ch} }}

\vspace*{1cm}

\centerline{\textsc{Abstract}}

\begin{center}
\parbox{10cm}{{ \small 
\hspace*{.1ex} %The vanishing of reduced $\ell^2$-cohomology for amenable groups can be traced to the work of Cheeger \& Gromov in \cite{CG}. Whether the same thing can be said for other $p \in ]1,\infty[$ still remains open. 
In this paper, it is shown that the reduced $\ell^p$-cohomology is trivial for a class of finitely generated amenable groups called transport amenable. These groups are those for which there exist a sequence of measures $\xi_n$ converging to a left-invariant mean and such that the transport cost between $\xi_n$ displaced by multiplication on the right by a fixed element and $\xi_n$ is bounded (uniformly in $n$). This class contains groups with controlled F{\o}lner sequence (such as polycylic groups) as well as some wreath products (such as $H' \wr H$ where $H$ is finitely generated Abelian and $H'$ is finitely generated amenable).
}}
\end{center}

%-------------------------------------------------------------------------
%-------------------------------------------------------------------------
\section{Introduction}\label{s-intro}
%-------------------------------------------------------------------------
%-------------------------------------------------------------------------

\setcounter{teo}{0}
\renewcommand{\theteo}{\thesection.\arabic{teo}}

% \subsection{Result}
%------------------

% \setcounter{teo}{1}

The study of $\ell^p$-cohomology for discrete spaces has been introduced by Gromov in \cite[\S{}8]{Gro} as an asymptotic invariant of groups (it turns out it is a quasi-isometry invariant, see below). 
The subject matter of this paper is the reduced $\ell^p$-cohomology in degree one of finitely generated groups. Let $\Gamma= (\Gver, E)$ be a Cayley graph (for a finite generating set), the reduced $\ell^p$-cohomology in degree one is the quotient 
\[
\ssl{\ell^p H}^1(\Gamma) := ( \ell^p(E) \cap \nabla \kk^\Gver ) / \srl{\nabla \ell^p(\Gver)}^{\ell^p(E)},
\]
where the gradient of a function on the vertices $g$ is defined by $\nabla g(\gamma,\gamma') = g(\gamma') - g(\gamma)$. 

When $G$ is a finitely generated group, this is isomorphic to the reduced cohomology of the left-regular representation on $\ell^p(G)$, see Puls' paper \cite{Puls-harm} or Martin \& Valette \cite{MV}. Another important result is that $\ell^p$-cohomology is an invariant of quasi-isometry:
\begin{tea}\emph{(see Élek \cite[\S{}3]{El-qi} or Pansu \cite{Pan-qi})}
If two graphs of bounded valency $\Gamma$ and $\Gamma'$ are quasi-isometric, then they have the same $\ell^p$-cohomology (in all degrees, reduced or not).
\end{tea}
The result is actually much more powerful, in the sense that it holds for a large category of measure metric spaces (see above mentioned references). For shorter proofs in more specific situations see Puls \cite[Lemma 6.1]{Puls-pharmbnd} or Bourdon \& Pajot \cite[Théorème 1.1]{BP}. As a consequence, if $G$ is a finitely generated group, the $\ell^p$-cohomology of any two Cayley graphs (for a finite generating set) are isomorphic. Thus, one may speak of the $\ell^p$-cohomology of a group without making reference to a Cayley graph.

The present paper gives a partial answers to a question (dating back at least to Gromov \cite[\S{}8.$A_1$.($A_2$), p.226]{Gro}):
\begin{ques}\label{laquestion}
Let $G$ be an amenable group, is it true that %for one (and hence all) Cayley graph $\Gamma$ and 
for all $1<p<\infty$, $\ssl{\ell^pH}^1(G)=\{0\}$?
\end{ques}
For $p=2$, the positive answer is a famous result of Cheeger \& Gromov \cite{CG} (see also Lück's book \cite{Luck}). The results presented here will cover other partial (positive) answers due to Kappos \cite{Kap}, Martin \& Valette \cite{MV} and Tessera \cite{Tes}.

The problem is reduced to a question of transport cost for measures tending to a left-invariant mean in $G$. %Here is the promised formulation in terms of transportation theory. 

\begin{dfn}
A finitely generated group $G$ is called transport amenable if there exists $S$ a finite generating set, a positive constant $K \in \rr$ and a sequence of finitely supported probability measure $\xi_n$ converging to a left-invariant mean such that, $\forall s \in S$ and $\forall \alpha$, the transport cost
\[
\trc(\rho_s \xi_n^\vee, \xi_n^\vee) = \inff{m \in M(\rho_s \xi_n^\vee, \xi_n^\vee)}  \int_{\Grou \times \Grou} d_\Gamma(x,y) \dd m(x,y) \leq K,
\]
where $d_\Gamma$ is the distance in the Cayley graph $\Gamma = \cay(G,S)$, $\xi^\vee(x) = \xi(x^{-1})$ (so that $\rho_s \xi^\vee = (\lambda_s \xi)^\vee$),  and $M(\xi,\phi)$ is the collection of finitely supported measures on $\Grou \times \Grou$ with marginals $\xi$ and $\phi$. 
\end{dfn}
\begin{teo}\label{tvantran-t}
Let $G$ be a finitely generated group. If $G$ is transport amenable then the degree one reduced $\ell^p$-cohomology of $G$ vanishes for all $p \in ]1,\infty[$. 
\end{teo}
The proof relies almost solely on classical functional analysis; the exception being a lemma from Holopainen \& Soardi \cite[Lemma4.4]{HS}.

This theorem extends a result of Tessera \cite[Theorem 2.2]{Tes} (when the $L^p$-representation is the left-regular representation, see theorem \ref{tconfol-t} in the present text) concerning groups with controlled F{\o}lner sequence (henceforth abbreviated by CF). Tessera's result implies these have trivial reduced degree one $\ell^p$-cohomology. The simplest example of a F{\o}lner sequence which is not CF is a sub-sequence of balls in a group of intermediate growth.

The exact range (in the class of amenable groups) of Tessera's result (and, consequently, of theorem \ref{tvantran-t}) is unclear, though Tessera's article \cite{Tes-iso} gives a temptingly complete description of such groups. This class contains groups of polynomial growth, polycyclic groups and wreath products $F \wr \zz$ where $F$ is a finite group.

However, theorem \ref{tvantran-t} applies to groups where it is known that \cite[Theorem 2.2]{Tes} does not apply. To see this, one must invoke the results of Erschler \cite{Ers} (or look again in \cite{Tes-iso}). Indeed, groups with CF have an isoperimetric profile in $\log v$, so that $\zz \wr \zz$ and $\zz_2 \wr \zz^2$ may not have CF. 

In examples \ref{ex-reverb} and \ref{ex-reverb2}, a small class of groups is shown to be transport amenable (though they are not CF, as explained in the previous paragraph). These groups are wreath products $H' \wr H$ where $H'$ is amenable (finite or infinite) and $H$ is infinite Abelian. In particular, this means that theorem \ref{tvantran-t} applies to $\zz \wr \zz$ and $\zz_2 \wr \zz^2$.

It was known that groups with infinitely many finite conjugacy classes (\eg an infinite centre, \eg groups of polynomial growth\footnote{Groups of polynomial growth are nilpotent by Gromov's famous result \cite{Gro-polyn}, and nilpotent groups possess infinitely many finite conjugacy classes.}) have trivial reduced $\ell^p$-cohomology in degree $1$ (see Kappos' preprint \cite[Theorem 6.4]{Kap}; the case of infinite centre is explained in Gromov \cite[\S{}8]{Gro} and detailed in Puls \cite[\S{}5]{Puls-prem} and Martin \& Valette's paper \cite[Theorem.(iii)]{MV}). Note that the same method used for theorem \ref{tvantran-t} may be applied to recover this vanishing result, see theorem \ref{tcentrinf-t}).

In \cite{moi-poiss} the author developed another method to show (among other results) that groups with trivial Poisson boundary (for a SRW on the Cayley graph, \ie Liouville) and superpolynomial growth have trivial reduced $\ell^p$-cohomology (see \tpoiss). Though all groups with CF have trivial Poisson boundary, groups like $\zz \wr \zz$ and $\zz_2 \wr \zz^3$ have non-trivial Poisson boundary\footnote{Wreath products $H' \wr H$ are also never finitely presented, unless $H' = \{1\}$ or $H$ is finite, see \cite{Baumslag-wrpasfinpres}.} the (see Kaimanovich \cite[Theorem 3.3]{Kaimano}). Thus \ref{tvantran-t} is not covered by \tpoiss.

The methods presented here may be of use in non-amenable groups, but apart for groups with infinitely many finite conjugacy classes, the author could not find any other case. They also apply to graphs, but the proper conditions on graphs are not so convenient to formulate (\eg quasi-transitive action by quasi-isometries).

When $1<p<\infty$, it is known (see Puls \cite[\S{}3]{Puls-harm} or Martin \& Valette \cite[\S{}3]{MV}) that the existence of non-constant $p$-harmonic function (\ie $h \in \D^p(\Gamma)$ with $\nabla^* \maza_{p,p'} \nabla h = 0$, where $\maza_{p,p'}$ is the Mazur map defined by $(\maza_{p,p'} f)(\gamma) = |f(\gamma)|^{p-2} f(\gamma)$) is equivalent to the non-vanishing of reduced cohomology in degree $1$. Other known consequences of the triviality of the reduced $\ell^p$-cohomology include the triviality of the $p$-capacity between finite sets and $\infty$ (see \cite{Yam} and \cite{Puls-pharmbnd}).

{\it Acknowledgments:}
The author is grateful to M.~de~la~Salle, P.~Pansu, J.~C.~Sikorav and R.~Tessera for comments and corrections.

%------------------------------------------
\section{Notations and preliminary definitions}
%------------------------------------------

Given a finitely generated group $G$ and a finite set $S$, the Cayley graph $\cay(G,S)$ is the graph whose vertices are the element of $G$ and $(\gamma,\gamma') \in E$ if $\exists s \in S$ such that $s^{-1} \gamma = \gamma'$. (Consequently, all Cayley graphs will be assumed to be of bounded valency.) In order for the resulting graph to have a symmetric edge set, $S$ is always going to be symmetric (\ie $s \in S \imp s^{-1} \in S$). Also, Cayley graphs are always going to be asumed connected (\ie $S$ is generating). 

The set of edges is thus thought of as a subset of $\Gver \times \Gver$. 
% The set of edges will be assumed symmetric (\ie $(x,y) \in E \imp (y,x) \in E$). 
Functions will take value in $\kk$, a complete normed field where classical theorems about Banach spaces hold, \eg $\rr$ or $\mathbb{C}$. 
% Furthermore, functions on $E$ will always be assumed to be anti-symmetric (\ie $f(x,y)= -f(y,x)$). 
Thus $\ell^p(G)$ is the Banach space of functions on the vertices which are $p$-summable, while $\ell^p(E)$ will be the subspace of %(anti-symmetric) 
functions on the edges which are $p$-summable. 

The gradient $\nabla:\kk^\Gver \to \kk^E$ is defined by $\nabla g(\gamma,\gamma') = g(\gamma') - g(\gamma)$, as the graphs are of bounded valency it is a bounded (and injective) operator from $\ell^p(G)$ into $\ell^p(E)$. Its kernel in $\kk^G$ is the space of constant functions. It is worthwhile to observe that the gradient is made of $ \{(\lambda_s - \Id) g\}_{s \in S}$ where $\lambda$ is the left-regular representation. As for the right-regular representation, it is a (injective) homomorphism from $G$ into $\mathrm{Aut}\big(\cay(G,S) \big)$, the automorphism group of the Cayley graph.

The Banach space of $p$-Dirichlet functions is the space of functions $f$ on $\Gver$ such that $\nabla f \in \ell^p(E)$. It will be denoted $\D^p(\Gamma)$. In order to introduce the $\D^p(\Gamma)$-norm on $\kk^\Gver$, it is necessary to choose a vertex $e$ (in a Cayley graph, it is convenient to choose the neutral element). This said $\|f\|_{\D^p(\Gamma)}^p = \|\nabla f\|_{\ell^p(E)}^p + |f(e)|^p$. By taking the primitive of the gradients, one may also prefer to define the reduced $\ell^p$-cohomology as 
\[
\ssl{\ell^p H}^1(\Gamma):= \D^p(\Gamma) / \srl{ \ell^p(\Gver) + \kk}^{\D^p}. 
\]

Lastly, $p'$ will denote the H\"older conjugate exponent of $p$, \ie $p' = p/(p-1)$ (with the usual convention that $1$ and $\infty$ are conjugate). 

A sequence of finite sets $\{F_n\}$ in $G$ will be called a left-F{\o}lner sequence (resp. right-F{\o}lner sequence) if
\[
\forall \gamma \in G, \frac{|\gamma F_n \Delta F_n|}{|F_n|} \tosp{n} 0, \quad \textrm{ (resp. } \forall \gamma \in G, \frac{|F_n \gamma \Delta F_n|}{|F_n|} \tosp{n} 0 \textrm{ )}.
\]
A finitely generated group is amenable if and only if it has a F{\o}lner sequence. Following Tessera \cite{Tes}, the F{\o}lner sequence will be called \emph{controlled}, if the above ratio is less than $K (\diam F_n)^{-1}$ for some constant $K$. 

Existence of a left- (resp. right-) F{\o}lner sequence is equivalent to the existence of an invariant mean, \ie a positive, linear, continuous, left- (resp. right-) translation invariant map $\mu: \ell^\infty(G) \to \rr$, often normalised so that $\mu(\un_G)=1$. This, in turn, implies the existence of a sequence of finitely supported positive probability measures $\xi_n$ satisfying $\|\lambda_\gamma \xi_n - \xi_n\|_{\ell^1(G)} \to 0$ (resp. $\|\rho_\gamma \xi_n - \xi_n\|_{\ell^1(G)} \to 0$). It is a standard trick to construct a bi-invariant (\ie left- \emph{and} right-) version of any of the previous items from a left- (resp. right-) invariant one. There are many possible references on this topic, \eg Paterson's \cite{amen} or Pier's \cite{amen2} book.

Throughout this text, $\delta_\gamma$ is the Dirac mass at $\gamma$: $\delta_\gamma(\eta) = 0 $ if $\eta \neq \gamma$ and $1$ if $\eta = \gamma$. The following convention for the (left-)convolution of function will be in use:
\[
\xi * f (\eta) = \sum_{\gamma \in G} \xi(\gamma) \lambda_\gamma f(\eta) = \sum_{\gamma \in G} \xi(\gamma) f(\gamma^{-1} \eta) 
\]
Also, since this section deals exclusively with Cayley graphs of a group $G$, the vertex set $\Gver$ is equal to (and thus replaced by) $\Grou$.

\setcounter{teo}{0}
\renewcommand{\theteo}{\thesection.\arabic{teo}}

\section{A criterion for convergence in $\D^p$}\label{ss-trique}
%-----------------------------------------------------------

Assume $\Gamma = \cay(G,S)$ is the Cayley graph of $G$ for the set $S$. The gradient $\nabla$ is made up of the partial gradients $\nabla_s:\kk^\Grou \to \kk^\Grou$ defined by 
\[
\nabla_s f = (\lambda_s - \Id) f = (\delta_s - \delta_e) * f. 
\]
Thus, if $f \in \D^p(\Gamma)$ then $\lambda_\gamma f$ represents the same class in $\ssl{\ell^p H}^1(\Gamma)$. Indeed, $\lambda_s f - f = \nabla_s f \in \ell^p(\Grou)$ so, writing $\gamma$ as a word in $S$, $\lambda_\gamma f$ is also in the same class. Furthermore, a convex combination of such functions will always represent the same class. Actually, if $\xi \in \kk(\Gamma)$ is of a function finite support with $\sum_{\gamma \in \Grou} \xi(\gamma) = 1$ then $[\xi * f] = [f]$.

Next recall that if a sequence $f_n \in \D^p$ converges point-wise (that is its value at $e$ and the values of the gradient on the edges) to $0$ and is bounded, then it weak$^*$-converges to $0$. This is relatively general fact. Indeed, assume $f_n \in \ell^p(\nn)$ is a bounded sequence which converges point-wise. Let $h$ be in the pre-dual, for any $\eps >0$, there is a finite set $F \subset \nn$ such that, if $h|_F$ is the restriction of $h$ to $F$, then $\|h - h|_F\| < \eps$. In particular, $|\pgen{f_n,h}| < |\pgen{f_n,h|_F}| + \|f_n\| \eps \leq |\pgen{f_n, h|_F}| + K \eps$.

Next, if $p \in ]1,\infty[$, weak$^*$-convergence of $f_n$ implies weak-convergence of $f_n$, which in turns implies norm-convergence of some sequence $h_k$ where each $h_k$ is a convex combination of the $f_n$ for $n \in \{1,2,\ldots ,k\}$ (a consequence of the Hahn-Banach theorem, see \cite[Theorem 3.13]{Rud}). Putting these two facts together one gets
\begin{lem}\label{tcondvan-l}
Let $f \in \D^p(\Gamma)$. If there exists a sequence $\xi_n$ of finitely supported functions such that:
\begin{enumerate}\renewcommand{\labelenumi}{{\normalfont \bf \arabic{enumi}.}}
\item $\sum_{\gamma \in \Grou} \xi_n = 1$;
\item $\forall s \in S, (\delta_s - \delta_e) * \xi_n * f$ converge point-wise to $0$; 
\item and $\exists K>0$ such that $\forall s \in S, \| (\delta_s - \delta_e) * \xi_n * f\|_{\ell^p(\Grou)} <K$,
\end{enumerate}
then $[f] = 0 \in \ssl{\ell^p H}^1(\Gamma)$.
\end{lem}
\begin{proof}
Indeed if {\bf 1} holds, then all the $\xi_n * f$ represent the same cohomology class. One may also add a constant function at every step, so that in addition to {\bf 2} these converges point-wise (at $e$). {\bf 3} implies that up to taking a convex combination, there is a sequence $y_n$ with $[y_n]=[f]$ (being convex combinations of $\xi_n * f$ and since $[\xi_n * f] = [f]$) and $y_n \to 0$ in the norm of $\D^p(\Gamma)$. This implies that the reduced class trivial. 
\end{proof}
This gives the following theorem, which may also be found in Kappos' preprint \cite{Kap} (where the conclusion extends to higher degrees if $G$ is of type $\FP_n$). %or Martin \& Valette \cite[Theorem 4.2]{MV} (the hypothesis there is either infinite centre in $G$ or a non-amenable normal subgroup $N$ with infinite centraliser $Z_G(N)$).
\begin{teo}\label{tcentrinf-t}
Let $G$ be a finitely generated group with infinitely many finite conjugacy classes. Then $\ssl{\ell^p H}^1(\Gamma)=0$ for any Cayley graph $\Gamma$ of $G$. 
\end{teo}
\begin{proof}
Take a sequence of $C_n \subset G$ of finite conjugacy class such that the distance to the identity tends to infinity and let $\xi_n = \un_{C_n}/|C_n|$ be the normalised characteristic function. Then, the condition of the above lemma are satisfied for any $f \in \D^p(\Gamma)$ as 
\begin{enumerate}\renewcommand{\labelenumi}{{\normalfont \bf \arabic{enumi}.}}
\item $\xi_n$ is positive and has $\|\xi_n\|_{\ell^1(G)} =1$;
\item Functions in $\ell^p(E)$ decay at infinity. Since  $(\delta_s -\delta_e) * \xi_n * f = \xi_n * (\delta_s -\delta_e) * f$ the point-wise convergence to $0$ follows;
\item Again $(\delta_s -\delta_e)* \xi_n * f = \xi_n * (\delta_s -\delta_e)*  f$ and since $\|\xi_n\|_{\ell^1(G)} = 1$ it has norm $1$ as a convolution on $\ell^p$, thus lemma \ref{tcondvan-l} may be applied with $K=1$.
\end{enumerate}
Since the above holds for any $f \in \D^p$, the $\ell^p$-cohomology ($1<p<\infty$) of groups with infinitely many finite conjugacy classes is trivial.
\end{proof}
As noted by Kappos in \cite{Kap}, the previous theorem covers nilpotent groups. For more amusing examples of groups with this property, see \cite{BCD}.

By a lemma of Holopainen \& Soardi \cite[Lemma 4.4]{HS} (and the equivalence between non-existence of non-constant $p$-harmonic maps and vanishing of the reduced $\ell^p$-cohomology in degree one due to Puls \cite[\S{}3]{Puls-harm}, see also Martin \& Valette \cite[\S{}3]{MV}), it is actually sufficient to show the vanishing of cohomology on bounded functions, \ie for $f \in \BD^p(\Gamma) := \ell^\infty(\Grou) \cap \D^p(\Gamma)$. This is quite useful, as one is tempted to take $\xi_n$ to be a sequence which tends to an invariant mean. 

\begin{lem}\label{tconvoper-l}
Let $G$ be an amenable group, $\Gamma= \cay(G,S)$ and $\xi_n$ is a sequence of measures tending to a left-invariant mean. If there is a $K>0$ such that 
\[
\forall s \in S, \quad \| (\delta_s - \delta_e) * \xi_n * \cdot \|_{\D^p(\Gamma) \to \ell^p(G)} \leq K
\]
then $\ssl{\ell^pH}^1(\Gamma)$=0.
\end{lem}
\begin{proof}
Indeed, assume $f \in \BD^p(\Gamma)$, then the first two points of lemma \ref{tcondvan-l} are automatically satisfied. Indeed, \textbf{1} is a consequence that $\xi_n$ is a probability measure and \textbf{2} follows from
\[
|(\delta_s - \delta_e) * \xi_n * f (\eta)| \leq \|\lambda_s \xi_n - \xi_n \|_{\ell^1(\Grou)} \|f\|_{\ell^\infty(\Grou)},
\]
as $\|\lambda_s \xi_n - \xi_n \|_{\ell^1(\Grou)} \to 0$. Finally, the hypothesis is a restatement of \textbf{3}. 
\end{proof}
In general, a computation gives:
\[ %\eqtag \label{eq-calcon}
\begin{array}{rl}
(\delta_s - \delta_e) * \xi_n * f (\eta) 
&= \displaystyle \sum_{\gamma \in s X_n} \xi_n(s^{-1} \gamma) \lambda_\gamma f(\eta) - \displaystyle \sum_{\gamma \in X_n} \xi_n(\gamma) \lambda_\gamma f(\eta)  \\
&= \displaystyle \sum_{\gamma \in X_n} \xi_n(\gamma) f(\gamma^{-1} s^{-1} \eta) - \displaystyle \sum_{\gamma \in X_n} \xi_n(\gamma) f(\gamma^{-1} \eta),
\end{array}
\]
where $X_n$ is the support of $\xi_n$. In order to express the above sum in terms of the gradient, it is necessary to express the difference as integrating the gradient on the paths. In the simpler case where $\xi_n = \un_{F_n} / |F_n|$ and $F_n$ is a left-F{\o}lner sequence, the above sum is a the sum of values of $f$ on $F_n^{-1} s^{-1} \eta \setminus F_n^{-1} \eta$ minus the values of $f$ on $ F_n^{-1} \eta \setminus F_n^{-1} s^{-1} \eta$, all this divided by $|F_n|$. So, from now on, the aim will be to bound the transport of a uniform mass on $F_n^{-1} s^{-1} \eta \setminus F_n^{-1} \eta$ towards a uniform mass on $ F_n^{-1} \eta \setminus F_n^{-1} s^{-1} \eta$ if the cost of transport is given by the gradient of $f$.

\section{Proof of the theorem \ref{tvantran-t}}\label{ss-preuve}
%-------------------------------------------------

The previous computation is an incentive to look at how one transports a probability measure to another.
\begin{dfn}
Let $\xi$ and $\phi$ be two finitely supported functions with \linebreak $\sum_{\gamma \in \Grou} \xi(\gamma) = \sum_{\gamma \in \Grou} \phi(\gamma)$. Fix a ``price'' function $f':E \to \kk$ on the edges (not necessarily positive). Let $\wp$ be the set of (finite) paths in the graph. 
% Fix a choice of paths $P: \Grou \times \Grou \to \wp$ such that $P(x,y)$ is a path from $x$ to $y$. 
Let $M_\wp(\xi,\phi)$ be the collection of (finitely supported) measures on $\wp$ such that the marginals are $\xi$ and $\phi$, \ie
\[
\sum_{P \in \wp, \textrm{source}(P)=x} m(P) = \xi(x) \qquad \textrm{and} \qquad \sum_{P \in \wp, \textrm{target}(P)=y} m(P) = \phi(y).
\]
Let $m \in M_\wp(\xi,\phi)$, %The transport cost $c_{f',m}: \Grou \times \Grou \to \rr_{\geq 0}$ is defined by $c_{f',P}(x,y) = \sum_{a \in P(x,y)} f'(a)$. 
$m$ is a transportation pattern or, for short, a pattern. Define the ``flattened'' measure $m^\flat$ on edges by the quantity of mass transported through that edge, \ie
\[
a \in E, \quad m^\flat(a) = \sum_{P \ni a} m(P),
\]
where the sum is to be made with multiplicities (if $a$ appears twice in $P$). The transportation cost of the measure $\xi$ to the measure $\phi$ according to the pattern $m$ and price $f'$ is
\[
\trc_{f',m}(\xi,\phi) =  \int_{E}  f'(a) \dd m^\flat(a).
\]
The infimal transportation cost of the measure $\xi$ to the measure $\phi$ is 
\[
\begin{array}{rll}
\trc(\xi,\phi) 
& := \displaystyle \inf_{m \in M_\wp(\xi,\phi)} \trc_{\un,m}(\xi,\phi) 
& = \displaystyle \inf_{m \in M_\wp(\xi,\phi)} \displaystyle  \int_{E} 1 \dd m^\flat(a) \\
& = \displaystyle \inf_{m' \in M_\Grou(\xi,\phi) } \displaystyle \int_{\Grou \times \Grou} d_\Gamma(x,y) \dd m'(x,y).
\end{array}
\]
where $\un$ is the function with constant value $1$, $M_\Grou(\xi,\phi)$ is the space of finitely supported measures on $\Grou \times \Grou$ with (usual) marginals $\xi$ and $\phi$ and $d_\Gamma(x,y)$ is the graph distance. \hfill $\Diamond$
\end{dfn}
Assume without loss of generality that we are dealing with probability measures. Notice that if $f' = \nabla f$, then for any pattern $m$,
\[
\begin{array}{rl}
\trc_{\nabla f,m}(\xi,\phi) 
  &= \displaystyle \int_{E} \nabla f(a) \dd m^\flat(a) \\
  &= \displaystyle \int_{\wp}  \sum_{a \in P} \nabla f(a) \dd m(P) \\
  &= \displaystyle \int_{\wp}  ( f(\text{target}(P)) - f(\text{source}(P)) ) \dd m(P) \\
%  &= \int_{\Grou \times \Grou} f(y) \dd m(x,y) - \int_{\Grou \times \Grou} f(x) \dd m(x,y) \\
  &= \displaystyle \int_{\Grou} f(y) \dd \phi(y) - \int_{\Grou} f(x) \dd \xi(x). \\
\end{array}
\]
Using this language one has that, for any pattern $m$, 
\[
(\delta_s - \delta_e) * \xi_n * f (\eta) = \trc_{\rho_\eta \nabla f,m}\big( (\lambda_s \xi_n)^\vee, \xi_n^\vee \big)
\]
where $\rho_\eta$ is the right regular representation and $\xi^\vee(\gamma) = \xi(\gamma^{-1})$ (so $(\lambda_s \xi)^\vee = \rho_s \xi^\vee$). 

\begin{proof}[Proof of theorem \ref{tvantran-t}]
The goal is to obtain the bound in lemma \ref{tconvoper-l}. By hypothesis, for every $n$, there is a pattern $m_n$ such that $\trc_{\un,m_n}(\rho_s \xi_n^\vee, \xi_n^\vee) \leq 2K$. For each $n$, this pattern will be used to get a bound on the map $T$ which sends a function $f' \in \ell^p(E)$ to $(Tf')(\eta) = \trc_{\rho_\eta f',m_n}\big( (\lambda_s \xi_n)^\vee, \xi_n^\vee \big)$ (which will be in $\ell^p(G)$). The bound will be obtained for $p=1$ and $p=\infty$, and deduced by Riesz-Thorin interpolation for intermediate $p$. In an attempt to alleviate notations the index $n$ will be omitted.

To get a $\ell^\infty$ bound, it suffices, to look at $\trc_{\un,m}(\rho_s \xi^\vee, \xi^\vee)$: right multiplication is an automorphism and replacing $f'$ by $\| f'\|_{\ell^\infty(E)}$ on each edges may only increase these sums. The $\ell^1$-bound is in appearance more intricate as one should not dispose so carelessly of the values of $f'$. Indeed,  if the weight function on the edges is $f' \in \ell^1(E)$ then the graph can be seen as a bounded (but not compact) metric space (\emph{via} $|f'|$).

The question can then be reduced to determining whether it is possible to solve (with $\xi = \rho_s \xi_n^\vee$ and $\phi = \xi_n^\vee$) the transport problem (and all its translates) so that $\forall a \in E, \sum_{\eta \in \Grou} m^\flat_{\rho_\eta f'; \xi,\phi}(a) < K \| f' \|_{\ell^1(E)}$ for any $f' \in \ell^1(E)$. This seems quite daunting. However, this bound is automatic if one has that $m^\flat_{\un; \xi , \phi} (E) \leq K$. Indeed, if one knows that the total quantity of mass is transported over all edges is bounded in one transport, then in all translations, any given edge will be used at most $|S|$ times this amount (the factor $|S|$ is necessary as the action of $G$ on non-oriented edges is not always faithful). In other words,
\[
\sum_{\eta \in \Grou} m^\flat_{\rho_\eta f'; \xi,\phi}(a) \leq |S| \trc_{\un} (\xi,\phi)  \| f' \|_{\ell^1(E)}. 
\]
Thus, using the Riesz-Thorin theorem, one gets a uniform bound on $\| (\delta_s - \delta_e) * \xi_n * \cdot \|_{\D^p(\Gamma) \to \ell^p(G)}$ for all $p \in [1,\infty]$ and, using lemma \ref{tconvoper-l}, shows that the only class in reduced cohomology is the trivial one. 
\end{proof}

Note that using $\xi_n = \un_{F_n} / |F_n|$ in theorem \ref{tvantran-t} (here $F_n$ are F{\o}lner sets), one sees immediately vanishing of cohomology for Abelian groups: there is a bijection between $F_n^{-1} s = s F_n^{-1}$ and $F_n^{-1}$ which sends every vertex to a neighbour using only one edge, so the above may be applied with $K=1$.

The hypothesis of theorem \ref{tvantran-t} are always satisfied when $G$ has a controlled F{\o}lner sequence, \ie a sequence of finite sets such that $\forall s \in S,$
\[
\frac{|F_n \Delta s F_n|}{|F_n|} \leq \frac{K}{ \diam F_n }.
\]
Since $F_n$ may be translated to be contained in the ball of radius $\diam F_n$. This gives back a special case (when the $L^p$-representation is the left-regular representation) of a result of Tessera \cite[Theorem 2.2]{Tes}:
\begin{teo}\label{tconfol-t}
If $G$ is a finitely generated group with a controlled F{\o}lner sequence, then its degree one reduced $\ell^p$-cohomology is trivial.
\end{teo}
\begin{proof}
If $\xi_n = \un_{F_n} /|F_n|$ one sees, up to translating $F_n$ into a ball of radius $\diam F_n$ around $e_G$, that the transport problem may be solved by picking any bijection from $F_n^{-1} s \setminus F_n^{-1}$ to $F_n^{-1} \setminus F_n^{-1} s$ and looking at the corresponding transport cost (at most $(2 \diam F_n + 2) |s F_n \Delta F_n| / |F_n|$).
\end{proof}

%----------------
\section{Examples}\label{ss-exemple}
%----------------

However, there are examples of groups which are not covered by Theorem \ref{tconfol-t} but where Theorem \ref{tvantran-t} applies. This is the case for many wreath products. The conventions for the wreath product (of two finitely generated groups) $H' \wr H$ will be as follows. $H'$ is the ``lamp state'' group and $H$ is the ``lamplighter position'' group. The generators are going to be of the form:  $S= \{e_H \to S_{H'}\} \times \{e_H \} \cup \{0\} \times S_H$ where $S_{H'}$ and $S_H$ are some set of generators for $H'$ and $H$ (respectively), $\{e_H \to S_{H'} \}$ is the set of functions supported on $e_H$ with value in $S_{H'}$ and $\{0\}$ is the trivial function (\ie sending all $H$ to $e_{H'}$). 

Multiplication on the left by $(e_H \to s_{H'}, e_H )$ changes the lamp state where the lamplighter is, whereas multiplication on the right changes the lamp state at the position $e_H$. Multiplication on the left by $(0, s_H)$ moves the lamplighter, whereas multiplication on the right moves the lamplighter together with all the lamps. 

\begin{exa}\label{ex-reverb}
Let $\Gamma$ be the Cayley graph of $H' \wr H$ where $H'$ is finite and $H$ is Abelian. 

A F{\o}lner sequence is given as follows: let $A_n$ be a controlled F{\o}lner sequence of diameter $d_n$ in $H$ and let $B_k$ be the functions $H \to H'$ which are supported on $A_k$ (\ie all the non-trivial lamp states supported by $A_k$). Then $F_n := B_n \times A_n$ is a (left- but not right-) F{\o}lner sequence. The transport problem is particularly easy to solve for $\xi_n = \un_{F_n} / |F_n|$, the normalised characteristic function of the F{\o}lner sets. 

Indeed, $r \in \{e_H \to S_{H'} \} \times \{e_H\}$ does not displace the F{\o}lner set at all. If $z \in \{0\} \times S_H$, then the problem is to displace elements from $F_n^{-1} z$ to $F_n^{-1}$ with a cost which is bounded by a constant multiple of $|F_n|$. 

Since the $F_n$ are particularly easy to describe and the $F_n^{-1}$ more ungainly, it turns out to be easier to write down the problem by inverting everything: namely displace $z^{-1} F_n$ to $F_n$, where the cost of displacing $\gamma$ to $\gamma'$ is the length of the element $\eta$ such that $\gamma \eta = \gamma'$.

Elements in  $z^{-1} F_n \setminus F_n$ are those where the lamplighter moved out of $A_n$. Since $H$ is assumed Abelian, for elements where the lamp states are all off on $A_n \setminus z^{-1} A_n$ it is possible to  displace all these element back to elements in $F_n$ by multiplying on the right by $z$. This will contribute to the final total cost (after division by $|F_n|$) for less than $\dfrac{|A_n \setminus z^{-1} A_n| }{ |A_n| c^{|A_n \setminus z^{-1}A_n|}}$.
% solving the transport problem on $H$ (which is true since $H$ is CF). So far the cost is bounded by $K_H$ (the constant from the transport problem on $H$). 

The aim is now to send elements with some non-trivial lamp states on $\Delta_z A_n := z A_n \setminus A_n$ to similar elements with non-trivial lamp states on $\Delta'_z A_n := A_n \setminus z A_n $. Let $\phi: A_n z \setminus A_n \to A_n \setminus A_n z$ be a bijection. Proceed as follows:
\begin{itemize}
\item If there are $i$ non-trivial lamp states, move the lamplighter (and the lamps, since the multiplication is on the right) so that $i$ lamps are successively at the identity and change their states to the trivial one, then turn on the $i$ corresponding (by $\phi$) lamps before coming back to a neighbouring position; this takes at most $2i(d_n+k)+d_n$ steps as $A_n$ is of diameter $d_n$ and the diameter of the ``lamp states group'' $F$ is $k$; 
\item For $i$ fixed, the number of such elements is $c^{|A_n|-|\Delta_z A_n|} (c-1)^i \binom{|\Delta_z A_n|}{i}$ where $c=|F|$;
\item So the total of used edges is at most (recall $\sum_{i=1}^{N} (c-1)^i i \binom{N}{i} = N c^{N-1}$)
\[
\begin{array}{l}
c^{|A_n|-|\Delta_z A_n|} \displaystyle \sum_{i=0}^{|\Delta_z A_n|} (c-1)^i \binom{|\Delta_z A_n|}{i} \big( 2i(d_n+k) + d_n ) \\
\qquad \leq c^{|A_n|-1} \bigg( 2 (d_n+k) |\Delta_z A_n| + d_n c \bigg);
\end{array}
\]
\item As $|F_n| = c^{|A_n|} |A_n|$, one gets a bound of $\dfrac{2 |\Delta_z A_n| (d_n + k) + d_n c}{|A_n| c}$.
\end{itemize}
Since $A_n$ may be chosen to be a controlled F{\o}lner sequence ($H$ is assumed Abelian), $d_n / |A_n| \leq 1$ and $k/c \leq 1$ is fixed, this is uniformly bounded by $1+\eps+2K/c$ where $\eps > 0$ is arbitrarily small and $K$ comes from the controlled F{\o}lner ratio. (Actually, if $H \neq \zz$, then the constant is $\eps+2K/c$.) Hence, theorem \ref{tvantran-t} applies.

The sequence $F_n$ is not itself a controlled F{\o}lner sequence (unless $H = \zz$): $\diam F_n \geq |A_n|$, $|F_n| = |A_n| k^{|A_n|}$ and $|z F_n \Delta F_n| \geq |\Delta_z A_n| k^{|A_n|}$. \hfill $\Diamond$
\end{exa}

The strategy may very well extend to more general wreath products. Considering $\xi_n = \un_{F_n} *\un_{F_n^{-1}} / |F_n|^2$ to get some sort of bi-CF sequence might also be a good idea. The parameters $k=\diam H'$ and $c=|H'|$ have been carefully brought along in the above example, the reason being:
%for $H' \wr H$ where $H'$ is finite and $H$ is a group having controlled left-F{\o}lner sequence but by 

\begin{exa}\label{ex-reverb2}
Let $H'$ and $H$ be infinite amenable groups with respective F{\o}lner sequences $A_n'$ and $A_n$ and respective generating sets $S_{H'}$ and $S_H$. Let $d_n=\diam A_n$ and $k_i=\diam A'_i$. Assume $H$ is Abelian and and that there is a sequence $i_n \in \nn$ such that 
\begin{itemize}
\item $\forall z \in S_{H'},  \limm{n \to \infty} \dfrac{2(d_n+k_{i_n}) |\Delta_z A'_{i_n}|}{|A_n| \, |A'_{i_n}|} \leq 2$.
%theorem \ref{tvantran-t} applies to $H'$ with $\xi_n$ the normalised characteristic function of $A'_n$; 
\item $\forall z \in S_{H},  \limm{n \to \infty} \dfrac{2(d_n+k_{i_n}) |\Delta_z A_n| + d_n |A'_{i_n}|}{|A_n| \, |A'_{i_n}|} \leq 2$.
\end{itemize}
Then theorem \ref{tvantran-t} applies to $H' \wr H$ (for the generating set $S= \{e_H \to S_{H'}\} \times \{e_H \} \cup \{0\} \times S_H$). 

To see this repeat the preceding example with $F_n$ being functions with support in $A_n$ and taking values in $A'_{i_n}$. (In doing so, $k$ becomes $k_i=\diam A'_i$ and $c$ becomes $c_i=|A'_i|$). The first condition comes up when solving the transport problem for the generators $(\{e_H \to h'\},e_H)$ (which no longer leave the set invariant). There are $|\Delta_z A'_{i_n}|$ which escape the F{\o}lner set, and each of them may be returned by a cost of at most $2(d_n+k_{i_n})$. The second is used when solving the transport problem for the generators $(e_{H'},h)$. 

The required sequence $i_n$ always exists. The first ratio always tends to $0$ if $i_n$ grows slowly enough, and, if $A_n$ is taken to be CF, the second ratio tends to either $1$ (when $H =\zz$) or $0$ (otherwise). \hfill $\Diamond$
\end{exa}

This means one may apply theorem \ref{tvantran-t} to $H' \wr \zz^k$ for any $H'$ amenable and $k >0$.

\end{document}